# DiffTester: Accelerating Unit Test Generation for Diffusion LLMs via Repetitive Pattern

### v2

@@ 4.2 Main Results
We conduct evaluations on the three benchmarks and two representative models described in Section 4.1. We use different batch sizes $n$, where $n$ ranges from 1 to 7, corresponding to the number of test cases generated per focal method. Figure 4 presents the average test coverage achieved under varying time budgets, while Table 1 reports the efficiency comparison between DiffTester and the baseline when generating the same number of test cases.
@@
Under the same time budget, DiffTester achieves higher test coverage. In pipeline of most software development, the time available for unit testing is often limited, which makes maximizing test coverage within a fixed time budget highly desirable. As shown in Figure 4, when the same amount of time is spent generating test cases, DiffTester consistently achieves substantially higher line coverage than the baseline. This demonstrates that the proposed DiffTester is highly competitive and enables dLLMs to achieve higher test coverage more quickly in UTG.
DiffTester does not compromise the maximum achievable test coverage. The highest test coverage attainable without time constraints is another important optimization target, particularly for scenarios such as vehicle control systems conrad2017testing; zhang2024visual. We find that DiffTester has little to no negative impact on the maximum test coverage and in some cases even improves it. For example, as shown in Figure 4, on the TestEval-Python benchmark with the Dream model, applying DiffTester leads to a slight increase in max test coverage.
@@
DiffTester can effectively improve the efficiency of dLLMs for UTG. Across nearly all settings, dLLMs achieve more than a 1.5× improvement in efficiency when equipped with DiffTester. For example, with a batch size $n=3$ on TestEval-C++ using DiffuCoder, DiffTester reduces the computational cost and decoding time from 1217 TFLOPs and 14.4s to 430 TFLOPs and 6.0s, respectively, while increasing throughput from 9.7 TPS to 23.8 TPS. These results demonstrate that DiffTester can substantially accelerate dLLMs in unit test generation.
@@
DiffTester generalizes well across different models and programming languages. Across various models and programming languages, DiffTester exhibits consistent acceleration trends and similar effects on coverage. Notably, the acceleration effect of DiffTester is more pronounced on TestEval-C++ compared to the other two benchmarks. For example, with a batch size of 5 using DiffuCoder, throughput improves by 1.55× on TestEval-Python and 1.71× on TestEval-Java, whereas on TestEval-C++ it increases by 2.45×. We attribute this to the fact that commonly used syntactic structures in C++ exhibit greater structural repetition than those in Python and Java, which makes it easier to exploit repetitive patterns for acceleration.
@@
Table 1: Efficiency comparison with and without DiffTester across different batch sizes $n$ on TestEval-Python, TestEval-C++, and TestEval-Java, with Dream and DiffuCoder.
@@ 5.1 Comparison with dLLM Acceleration for General Task
We observe that several contemporary works have also explored training-free approaches to accelerate dLLMs wei2025accelerating; ben2025accelerated; israel2025accelerating. However, these approaches typically focus on general-purpose tasks, whereas DiffTester is designed specifically for unit test generation (UTG), a domain where efficiency is particularly critical. To highlight the difference, we compare DiffTester with a representative approach, EB-Sampler ben2025accelerated. EB-Sampler accelerates dLLM inference through an Entropy-Bounded unmasking procedure, which dynamically unmasks multiple tokens under a predefined approximate error tolerance.
We conduct experiments on the TestEval-Python, testing batch sizes of $n=3,5,7$, and report both average decoding time and line coverage in Table 2. ❶ Compared with EB-Sampler, DiffTester achieves more effective acceleration in generating unit tests. For example, when $n=7$, DiffTester reduces the average decoding time to 17.6%, whereas EB-Sampler reduces it only to 19.8%. ❷ In addition, the quality of the test cases generated by DiffTester is substantially higher. With DiffTester, the line coverage remains almost identical to the baseline without any acceleration, while EB-Sampler leads to a significant drop in coverage. We attribute the superior performance of DiffTester in terms of both efficiency and quality to its ability to exploit repetitive structural patterns that are inherent to the UTG task. This task-specific insight allows DiffTester to accelerate generation without compromising the quality of the resulting test cases.
Table 2: Decoding time and line coverage on TestEval-Python with batch sizes $n=3,5,7$, comparing the baseline DiffuCoder, EB-Sampler, and DiffTester.
@@
+ DiffTester 7.8 12.6 17.6 92 93 94
@@ 5.2 Case Study
Figure 5: Case study on Dream illustrating the decoding process with and without DiffTester.
We present a concrete case in Figure 5 to further illustrate the impact of DiffTester. For the same focal method, we compare the decoding process of Dream with and without DiffTester. The case clearly shows that DiffTester substantially accelerates UTG: within 21 decoding steps, the version with DiffTester has already generated nearly complete and valid test cases, whereas the version without DiffTester still fails to produce a proper test case and contains numerous syntax errors.
5.3 Distribution of Inference Steps with DiffTester
In this section, we analyze the distribution of the number of inference steps after applying DiffTester. The experiments are conducted on DiffuCoder across three benchmarks, with the predefined generation length set to 128. Under the default remasking strategy, the model decodes two tokens at each step, which results in 64 inference steps without DiffTester. The distribution of inference steps after applying DiffTester is shown in Figure 6. We observe that the number of steps follows an approximately normal distribution, with most cases concentrated at values significantly smaller than 64. Moreover, the number of inference steps adapts to the characteristics of each focal method, making the approach well-suited for diverse testing scenarios.
Figure 6: Distribution of the number of inference steps after applying DiffTester on DiffuCoder. The red dashed line indicates the original fixed step number without acceleration.
6 Related Work
6.1 Unit Test Generation
Unit tests UT1; UT2 are essential for verifying the behavior of program units and serve as a foundational mechanism for early fault detection and prevention in software development. However, writing unit tests manually is highly labor-intensive, which has motivated extensive research on automated unit test generation dakhel2024effective; aniche2022effective. Traditional techniques primarily rely on search-based search_base; search-based2, random-based random_based, constraint-based constrained_based, and symbolic execution-based approaches symbolic_UT_1, with widely used tools such as EvoSuite EvoSuite, Pynguin Pynguin, and KLEE KLEE. Nevertheless, these traditional approaches often suffer from limitations such as low coverage time-consuming1.
With the rise of large language models (LLMs) achiam2023gpt, recent advances have increasingly explored their application to automated test case generation. Representative approaches include ChatTester ChatTester, TestPilot TestPilot, and ChatUnitTest ChatUnitTest. Although LLMs have significantly improved test coverage, existing LLMs used for unit test generation almost exclusively adopt a token-by-token generation paradigm zhang2025beyond, which constrains the efficiency of the unit test generation.
6.2 Diffusion Large Language Model
Diffusion language models have recently become a focal point in AI research sahoo2024simple; discreteVLA and can be categorized into continuous and discrete formulations li2025survey. Empirical evidence zhao2025d suggests that discrete diffusion language models scale more effectively to larger model sizes, which has led to the development of diffusion large language models such as LLaDA llada and Gemini Diffusion gemini_diffusion. Within this line of work, MMaDA yang2025mmada and LLaDA-V you2025llada extend dLLMs to the multimodal domain; TreeDiff zeng2025treediff explores integrating dLLMs with the structural characteristics of code; DLLM-Cache dLLM-Cache and Sparse-dLLM song2025sparse propose inference acceleration strategies based on KV-cache mechanisms; DiffuCoder DiffuCoder and LLaDA-1.5 zhu2025llada15variancereducedpreference investigate reinforcement learning tailored to dLLMs.
@@ 7 Conclusion
In this paper, we present an in-depth analysis of the pervasive repetitive patterns in unit test generation and the characteristics of dLLMs’ inference. Building on this analysis, we propose DiffTester, the first acceleration framework specifically designed for dLLMs in unit test generation. The central idea of DiffTester is to extract repetitive patterns in intermediate unit test cases and leverage them to decode a larger number of tokens per inference step. Extensive experimental results demonstrate that DiffTester can substantially accelerate dLLMs in unit test generation while preserving test coverage, and that it generalizes effectively across different dLLMs and programming languages. We believe that this work provides a practical and scalable solution for improving the efficiency of unit test generation in software development.
8 Reproducibility statement
To ensure the reproducibility of our results, we have made all relevant code and datasets publicly available at https://github.com/wellbeingyang/DLM4UTG-open. The repository includes detailed instructions for setting up the environment, running experiments, and reproducing all results presented in the paper. We encourage the community to use these resources to verify our findings and to facilitate further research in this area.
@@ Appendix A LLM Usage Statement
Appendix B Ethics Statement
@@ C.1.1 Models
We primarily conducted experiments using the dLLMs DiffuCoder-7B-cpGRPO (DiffuCoder) (DiffuCoder in this paper for short) and Dream-v0-Instruct-7B dream (Dream in this paper for short).
@@ C.1.3 Hyperparameters
In our experiments with the EB-Sampler method, we adopted the parameters recommended in the original paper ben2025accelerated. We used confidence as the error proxy function. And we used $\gamma=0.1$.
@@ C.1.4 Benchmark
To evaluate the performance of our method, we experiment on the TestEval benchmark testeval. TestEval is a benchmark for test case generation with LLMs. It includes 210 Python programs from an online programming platform, LeetCode. It has three metrics: overall coverage, targeted line/branch coverage, and targeted path coverage.
@@ C.2 More Experiments
C.2.1 Quality vs Speed
@@
Figure 7: Syntactic correctness of the code generated with different generation speed.
@@
C.2.2 Ablation Study on Threshold
@@
C.2.3 Ablation Study on Step Size
@@
C.2.4 Ablation Study on [PAD] Token
@@
Results are shown in Figure 11. The y-axis represents the average coverage divided by the average time or average computational cost, while the x-axis denotes the number of test cases generated simultaneously. The higher y-value means better performance. The curve labeled baseline means no acceleration method applied. The curve labeled +pad acceleration means applying this trick on the baseline, while the curve labeled +DiffTester means applying both this trick and our method.
@@ C.3 Prompts
You are a professional who writes {language} testcases. You always
respond without any natural language descriptions. Especially,
your answer should contain only one testcase.
@@
Please write a test method for the function ’{func_name}’ given
the following program under test and function description. Your
answer should only contain one test input.
@@
Function description for ’{func_name}’:
@@
Please write a test method for the function ’{func_name}’ given
the following program under test and function description. Your
answer should only contain one test input.
@@
Function description for ’{func_name}’:
@@
You can directly use ‘assertEquals‘ function for assertion. Your
testcase should be formatted as:
@@
Please write a testcase for the function ’{func_name}’ given the
following program under test and function description. Your answer
should only contain one test input.
@@
Function description for ’{func_name}’:
@@
You can directly use ‘assert‘ function for assertion. Your

### v3

@@ 4.2 Main Results
We conduct evaluations on the three benchmarks and two representative models described in Section 4.1. We use different batch sizes $n$, where $n$ ranges from 1 to 7, corresponding to the number of test cases generated per focal method. Figure 3 presents the average test coverage achieved under varying time budgets, while Table 1 reports the efficiency comparison between DiffuTester and the baseline when generating the same number of test cases.
@@
DiffuTester can effectively reduce the time needed to achieve a certain coverage rate. In pipeline of most software development, the time available for unit testing is often limited, which makes achieving a certain coverage as fast as possible highly desirable. As shown in Figure 3 and Table 1, to achieve the same coverage, DiffuTester can speedup up to $2\times-3\times$ compared with the baseline. For example, with a batch size $n=3$ on TestEval-C++ using DiffuCoder, DiffuTester reduces the decoding time from 14.4s to 6.0s, while increasing throughput from 9.7 TPS to 23.8 TPS. This demonstrates that the proposed DiffuTester is highly competitive and enables dLLMs to achieve higher test coverage more quickly in UTG.
@@
DiffuTester can effectively reduce the computational cost needed to generate test cases. Across nearly all settings, dLLMs achieve more than a $1.6\times$ improvement in computational cost when equipped with DiffuTester. For example, with a batch size $n=3$ on TestEval-C++ using DiffuCoder, DiffuTester reduces the computational cost from 1217 TFLOPs to 430 TFLOPs. These results demonstrate that DiffuTester can substantially save computational resources.
DiffuTester does not compromise the maximum achievable test coverage. The highest test coverage attainable without time constraints is another important optimization target, particularly for scenarios such as vehicle control systems Conrad and Fey (2017); Zhang et al. (2024). We find that DiffuTester has little to no negative impact on the maximum test coverage and in some cases even improves it. For example, as shown in Figure 3, on the TestEval-Python benchmark with the Dream model, applying DiffuTester leads to a slight increase in max test coverage.
@@
DiffuTester generalizes well across different models and programming languages. Across various models and programming languages, DiffuTester exhibits consistent acceleration trends and similar effects on coverage. Notably, the acceleration effect of DiffuTester is more pronounced on TestEval-C++ compared to the other two benchmarks. For example, with a batch size of 5 using DiffuCoder, throughput improves by 1.55× on TestEval-Python and 1.71× on TestEval-Java, whereas on TestEval-C++ it increases by 2.45×. We attribute this to the fact that commonly used syntactic structures in C++ exhibit greater structural similarity than those in Python and Java, which makes it easier to exploit structural patterns for acceleration.
Comparison with an AR baseline: We also conducted an experiment comparing dLLM equipped with DiffuTester with an auto-regressive model for reference. Results show that dLLMs equipped with DiffuTester require less time and computational cost to achieve the same coverage in most cases compared to auto-regressive models of similar scale. The details of the experiment are provided in Appendix C.2.1.
@@
Table 1: Efficiency comparison with and without DiffuTester across different batch sizes $n$ on TestEval-Python, TestEval-C++, and TestEval-Java, with Dream and DiffuCoder.
@@ 5.1 Comparison with dLLM Acceleration for General Task
Most training-free approaches to accelerate dLLMs can be classified into two categories: cache-based and sampling techniques. DiffuTester also focuses on optimizing the sampling process, making it compatible with cache-based techniques. Therefore, we only compare DiffuTester with some acceleration approaches of the same category.
Several contemporary works have also explored training-free approaches to accelerate dLLMs via advanced sampling strategies Wei et al. (2025); Ben-Hamu et al. (2025); Israel et al. (2025). However, such approaches typically target general-purpose tasks, rendering them suboptimal for UTG. To highlight the difference, we compare DiffuTester with two representative approaches, EB-Sampler Ben-Hamu et al. (2025) and SlowFast Sampling Wei et al. (2025).
EB-Sampler accelerates dLLM inference through an Entropy-Bounded unmasking procedure, which dynamically unmasks multiple tokens under a predefined approximate error tolerance. We conduct experiments on the TestEval-Python, testing batch sizes of $n=3,5,7$, and report both average decoding time and line coverage in Table 2.
❶Compared with EB-Sampler, DiffuTester achieves more effective acceleration in generating unit tests. For example, when $n=7$, DiffuTester reduces the average decoding time to 17.6s, whereas EB-Sampler reduces it only to 19.8s.
❷In addition, the quality of the test cases generated by DiffuTester is substantially higher. With DiffuTester, the line coverage remains almost identical to the baseline without any acceleration, while EB-Sampler leads to a significant drop in coverage.
Table 2: Decoding time and line coverage on TestEval-Python with batch sizes $n=3,5,7$, comparing the baseline DiffuCoder, EB-Sampler, and DiffuTester.
@@
+ DiffuTester 7.8 12.6 17.6 92 93 94
SlowFast Sampling is a novel dynamic sampling strategy that adaptively alternates between exploratory and accelerated decoding stages. We conduct experiments on the TestEval-Python, testing batch sizes of $n=7,10$ (for SlowFast Sampling) and report the average decoding time and line coverage in Table 3.
❶Although SlowFast Sampling is the fastest at the same testing batch size, it substantially degrades generation quality, achieving only 77% line coverage even when $n=10$, underperforming even the baseline configuration with $n=3$.
❷Moreover, SlowFast Sampling attains identical line coverage at $n=7$ and $n=10$, indicating that its peak coverage (77%) is significantly reduced by the acceleration strategy. In contrast, DiffuTester accelerates generation while preserving generation quality.
Table 3: Decoding time and line coverage on TestEval-Python, comparing the baseline DiffuCoder, SlowFast Sampling, and DiffuTester.
Method Decoding Time (s) Line Coverage (%)
$n=3$ $n=7$ $n=10$ $n=3$ $n=7$ $n=10$
DiffuCoder 12.2 26.8 36.1 91 94 94
+ SlowFast Sampling 2.4 5.4 7.7 73 77 77
+ DiffuTester 7.8 17.6 24.5 92 94 94
Compared to the two acceleration methods discussed earlier, DiffuTester can achieve good speedup without sacrificing generation quality. We attribute the superior performance of DiffuTester in terms of both efficiency and quality to its ability to exploit structural patterns that are inherent to the UTG task. This task-specific insight allows DiffuTester to accelerate generation without compromising the quality of the resulting test cases.
5.2 Applying DiffuTester to Other Code Generation Tasks
We also evaluate our method on other code generation tasks. Although these tasks may not exhibit repeated patterns as prominently as UTG, when multiple samples are generated, repeated patterns across samples can still be captured by DiffuTester and exploited to accelerate generation. For general code generation, we use the Python, C++ and Java subsets of the HumanEval-X benchmark Zheng et al. (2023). For code translation, we adopt datasets from UniTrans Yang et al. (2024). The experimental results are summarized in Table 4.
Table 4: Decoding time (s) and pass@1 (%) on other code generation tasks using DiffuCoder.
General Code Genration Python C++ Java
Time Pass@1 Time Pass@1 Time Pass@1
DiffuCoder 5.68 50 5.73 11 6.32 11
+ DiffuTester 3.87 49 4.21 13 5.36 11
Code Translation Python to C++ C++ to Java Java to Python
Time Pass@1 Time Pass@1 Time Pass@1
DiffuCoder 11.75 42.9 11.97 39.0 12.29 95.8
+ DiffuTester 10.34 46.2 10.06 40.1 10.00 97.5
Code Translation C++ to Python Java to C++ Python to Java
Time Pass@1 Time Pass@1 Time Pass@1
DiffuCoder 11.98 96.4 12.31 57.0 11.76 35.2
+ DiffuTester 9.59 95.5 10.57 56.1 10.26 36.2
We conducte the experiments using DiffuCoder to generate 4 samples each prompt and report the average decoding time and pass@1111Average pass rate of 4 samples. DiffuTester achieves a modest acceleration over the baseline. The acceleration is less pronounced than in UTG because repeated structures in these code generation tasks are not as salient as those in UTG. Nevertheless, DiffuTester generally does not degrade generation quality and, in some cases, even yields higher-quality outputs.
@@ 5.3 Case Study
Figure 4: Case study on Dream illustrating the decoding process with and without DiffuTester.
We present a concrete case in Figure 4 to further illustrate the impact of DiffuTester. For the same focal method, we compare the decoding process of Dream with and without DiffuTester. The case clearly shows that DiffuTester substantially accelerates UTG: within 21 decoding steps, the version with DiffuTester has already generated nearly complete and valid test cases, whereas the version without DiffuTester still fails to produce a proper test case and contains numerous syntax errors.
5.4 Distribution of Inference Steps with DiffuTester
Figure 5: Distribution of the number of inference steps after applying DiffuTester on DiffuCoder. The red dashed line indicates the original fixed step number without acceleration.
In this section, we analyze the distribution of the number of inference steps after applying DiffuTester. The experiments are conducted on DiffuCoder across three benchmarks, with the predefined generation length set to 128. Under the default remasking strategy, the model decodes two tokens at each step, which results in 64 inference steps without DiffuTester. The distribution of inference steps after applying DiffuTester is shown in Figure 5. We observe that the number of steps follows an approximately normal distribution, with most cases concentrated at values significantly smaller than 64. Moreover, the number of inference steps adapts to the characteristics of each focal method, making the approach well-suited for diverse testing scenarios.
@@ 6 Conclusion
In this paper, we present an in-depth analysis of the pervasive structural patterns in unit test generation and the characteristics of dLLMs’ inference. Building on this analysis, we propose DiffuTester, the first acceleration framework specifically designed for dLLMs in unit test generation. The central idea of DiffuTester is to extract structural patterns in intermediate unit test cases and leverage them to decode a larger number of tokens per inference step. Extensive experimental results demonstrate that DiffuTester can substantially accelerate dLLMs in unit test generation while preserving test coverage, and that it generalizes effectively across different dLLMs and programming languages. We believe that this work provides a practical and scalable solution for improving the efficiency of unit test generation in software development.
7 Limitations
Due to limited time and computational resources, we did not evaluate our method on additional models or code generation tasks. However, we believe that our approach is generally applicable to other dLLMs and code generation tasks.
8 Ethical Considerations
@@ Appendix A LLM Usage Statement
Appendix B Reproducibility statement
To ensure the reproducibility of our results, we have made all relevant code and datasets publicly available at https://github.com/TsinghuaISE/DiffuTester. The repository includes detailed instructions for setting up the environment, running experiments, and reproducing all results presented in the paper. We encourage the community to use these resources to verify our findings and to facilitate further research in this area.
@@ C.1.1 Models
We primarily conducted experiments using the dLLMs DiffuCoder-7B-cpGRPO (Gong et al., 2025) (DiffuCoder in this paper for short) and Dream-v0-Instruct-7B Ye et al. (2025) (Dream in this paper for short).
@@ C.1.3 Hyperparameters
In our experiments with the EB-Sampler method, we adopted the parameters recommended in the original paper Ben-Hamu et al. (2025). We used confidence as the error proxy function. And we used $\gamma=0.1$.
In our experiments with the SlowFast Sampling Wei et al. (2025) method, we adopted default parameters in the implementation by the authors. [^5] https://github.com/LiangrunFlora/Slow-Fast-Sampling/blob/main/slow-fast-sampling/sampling_utils/dream_generation_utils.py
@@ C.1.4 Benchmark
To evaluate the performance of our method, we experiment on the TestEval benchmark Wang et al. (2025). TestEval is a benchmark for test case generation with LLMs. It includes 210 Python programs from an online programming platform, LeetCode. It has three metrics: overall coverage, targeted line/branch coverage, and targeted path coverage.
@@ C.2 More Experiments
C.2.1 Comparison with AR model
In this section, we present a comparison between our method and an AR baseline model.
Figure 6: Dream + DiffuTester vs Qwen-2.5-7B-Instruct
From Figure 6, we can see that with DiffuTester, dLLMs require substantially less time and computational cost to achieve the same coverage in most cases compared to auto-regressive models of similar scale.
In this experiment, the AR baseline employs a strictly vanilla decoding procedure—without any acceleration techniques (e.g., speculative decoding Leviathan et al. (2023); Chen et al. (2023)) or quality-enhancing strategies (e.g., codamosa Lemieux et al. (2023)). The reason for not applying the acceleration techniques is that we primarily aim to explore the fundamental advantages of dLLMs over AR models at the foundation model level. Applying additional acceleration strategies to AR models would compromise the fairness of comparison, and moreover, these techniques could also be applied to dLLMs in the future. The reason for not applying the quality-enhancing strategies is that our goal is to accelerate UTG rather than to produce higher-quality unit tests. Such quality-enhancing techniques can likewise be incorporated to improve the test cases generated by dLLMs.
C.2.2 More Coverage Metrics
In this section, we present some experiment results using branch coverage.
Figure 7: Comparison of line/branch coverage with and without DiffuTester at equal decoding time.
From Figure 7, we can see that using branch coverage as a metric yields results that are highly consistent with those of line coverage. That is the reason why we only reported line coverage in the main results.
C.2.3 Quality vs Speed
@@
Figure 8: Syntactic correctness of the code generated with different generation speeds.
@@
C.2.4 Ablation Study on Threshold
@@
C.2.5 Ablation Study on Step Size
@@
C.2.6 Ablation Study on [PAD] Token
@@
Results are shown in Figure 12. The y-axis represents the average coverage divided by the average time or average computational cost, while the x-axis denotes the number of test cases generated simultaneously. The higher y-value means better performance. The curve labeled baseline means no acceleration method applied. The curve labeled +pad acceleration means applying this trick on the baseline, while the curve labeled +DiffuTester means applying both this trick and our method.
@@ C.3 Prompts
You are a professional who writes
{language} testcases. You always
respond without any natural
language descriptions.
Especially, your answer should
contain only one testcase.
@@
Please write a test method for
the function ’{func_name}’ given
the following program under test
and function description. Your
answer should only contain one
test input.
@@
Function description for
’{func_name}’:
@@
Please write a test method for
the function ’{func_name}’ given
the following program under test
and function description. Your
answer should only contain one
test input.
@@
Function description for
’{func_name}’:
@@
You can directly use
‘assertEquals‘ function for
assertion. Your testcase should
be formatted as:
@@
Please write a test method for
the function ’{func_name}’ given
the following program under test
and function description. Your
answer should only contain one
test input.
@@
Function description for
’{func_name}’:
@@
You can directly use ‘assert‘
function for assertion. Your